Sara Sharp, a 30-year-old insurance broker, decided to start a retirement plan. She figures that her income for the next 20 years will be sufficient to deposit $300 at the end of each month into her retirement plan. She will then let the money sit for another 10 years until she is 60 years old. If Sarah’s retirement plan earns 5.75% compounded monthly, what amount will she have when she turns 60? and How much interest did she accumulate on her money over the 30 years?
Answer by addingup(3677)

The first 20 years:
Future Value = 300[[[{1+(0.0575/12)}^(12*20)]-1]/(0.0575/12)]
= 300[[{(1.00479)^240}-1]/0.00479]
= 300(2.1483/0.00479)
= 300(448.49690) = 134,549.07 this is what she has at the end of 20 years. And she lets it sit for another 10 year earning 5.75% compounded monthly:
Future value = 134,549.07{(1+(0.0575/12))^(12*10)}
= 134,549.07(1.774) = 238,735.48 this is how much she will have at the end of 30 years